# GAIA: A Global, Multi-modal, Multi-scale Vision-Language Dataset for Remote Sensing Image Analysis

### v1

@@ Abstract
The continuous operation of Earth-orbiting satellites generates vast and ever-growing archives of Remote Sensing (RS) images. Natural language presents an intuitive interface for accessing, querying, and interpreting the data from such archives. However, existing Vision-Language Models (VLMs) are predominantly trained on web-scraped, noisy image-text data, exhibiting limited exposure to the specialized domain of RS. This deficiency results in poor performance on RS-specific tasks, as commonly used datasets often lack detailed, scientifically accurate textual descriptions and instead emphasize solely on attributes like date and location. To bridge this critical gap, we introduce GAIA, a novel dataset designed for multi-scale, multi-sensor, and multi-modal RS image analysis. GAIA comprises of 205,150 meticulously curated RS image-text pairs, representing a diverse range of RS modalities associated to different spatial resolutions. Unlike existing vision-language datasets in RS, GAIA specifically focuses on capturing a diverse range of RS applications, providing unique information about environmental changes, natural disasters, and various other dynamic phenomena. The dataset provides a spatially and temporally balanced distribution, spanning across the globe, covering the last 25 years with a balanced temporal distribution of observations. GAIA’s construction involved a two-stage process: (1) targeted web-scraping of images and accompanying text from reputable RS-related sources, and (2) generation of five high-quality, scientifically grounded synthetic captions for each image using carefully crafted prompts that leverage the advanced vision-language capabilities of GPT-4o. We also release an automated processing framework developed for this purpose, enabling the broader research community to generate captions for RS images using the web-crawled image-text data. Our extensive experiments, including fine-tuning of CLIP and BLIP2 models, demonstrate that GAIA significantly improves performance on RS image classification, cross-modal retrieval and image captioning tasks, proving its value as a crucial resource for advancing the field.
Index Terms:
Vision-language dataset, vision-language model, representation learning, remote sensing.
@@ I Introduction
The advancements in Vision-Language Models (VLMs) have drawn significant attention in the RS field, leading to studies across different applications. While early attempts have shown the potential of applying VLMs to RS, it is still an emerging field with many unsolved challenges [5]. A challenging issue that impedes the development of VLMs in RS is the lack of large-scale aligned image-text datasets. Existing RSI datasets mostly focus on visual recognition tasks and do not provide language annotations. While there have been some efforts to create RS image-text paired datasets, their scale, quality, and diversity still leave room for improvement.
@@
To address these challenges, we release GAIA, a dataset of approximately 41,030 RS images, each accompanied by five diverse and informative captions, resulting in 205,150 image-text pairs. Specifically, less that 3% of GAIA’s images are found in compared web-crawled datasets. Moreover, we provide extensive benchmarks for RS image classification and cross-modal retrieval tasks by fine-tuning CLIP on the GAIA dataset, and we publish the pre-trained model weights to facilitate further research. Complementing the dataset, we introduce a novel automated processing framework for generating RS image descriptions by leveraging web-crawled data from reputable sources and the advanced language generation capabilities of ChatGPT. Our work aims to bridge the gap between the vast potential of VLMs and the unique challenges of the RS domain, accelerating the development of foundation models that can unlock new insights from Earth observation data. We make our dataset, automated processing framework and fine-tuned model weights publicly available on our project’s GitHub repository [^2] https://github.com/Orion-AI-Lab/GAIA.
@@
    We introduce GAIA, a carefully curated RS dataset with 41,030 images along 5 synthetic captions each (205,150 image-text pairs) featuring global, multi-modal, multi-scale coverage spanning across the last 25 years. It exhibits minimal overlap ($<3\%$) with existing datasets, enabling diverse Earth event analysis across multiple spatial resolutions and modalities.
@@ II Related work
Vision-Language Models (VLMs) have emerged as powerful tools for understanding and generating connections between visual and textual information. Pioneering works like CLIP [11] demonstrated the effectiveness of contrastive learning in creating aligned visual-semantic embeddings, while more recent models such as LLaVA [12, 13] have expanded these capabilities to include sophisticated visual reasoning and open-ended dialogue. Modern VLMs can perform a wide range of tasks, including image captioning, visual question answering, and cross-modal retrieval without task-specific fine-tuning. The success of these models stems from their ability to learn generalizable visual-semantic representations through self-supervised learning on vast amounts of image-text data, typically utilizing web-scale datasets containing billions of image-text pairs. This approach has proven particularly effective at capturing nuanced relationships between visual content and natural language descriptions, enabling zero-shot transfer to novel tasks and domains.
@@ II-A Have existing VLMs seen Remote Sensing imagery?
Several research efforts have focused on extracting RS subsets from web-crawled large-scale general-purpose datasets, extensively used within the computer vision domain. As an example, Czerkawski and Francis developed LAION-EO [6], leveraging CloudSEN12 [15] as an ”anchor dataset” of Sentinel-2 images to retrieve similar images from LAION-5B. Their multi-stage filtering process, involving nearest neighbor search and refined CLIP-based similarity thresholds, yielded 24k samples in the prototype version and an expanded set of 113k in version 1. Similarly, Zhang et al. [16] identified LAION-RS, a subset of 726k RS images, i.e. a mere 0.03% of LAION-2B, using a binary classifier. Zhang et. al. [17] introduced RS5M, a dataset of 5 million RS image-text pairs. A core component of RS5M, the PUB11 subset containing over 3 million pairs, was created by filtering 11 large-scale public datasets using keyword-based filtering, de-duplication, and a VLM-based filter. The 11 filtered large-scale public datasets, include LAION2B-en [3], LAION400M [18], LAIONCOCO [19], COYO700M [20], CC3M [21], CC12M [22], YFCC15M [23], WIT [24], Redcaps [25], SBU [26], and Visual Genome [27]. These filtering efforts offer critical insights into the current state of VLMs. First, the relatively small size of the extracted subsets, such as the minuscule 0.03% of LAION-2B forming LAION-RS, demonstrates the limited and incidental exposure of VLMs to RS data. Second, as shown in Fig. 1, web-scraped datasets like LAION-EO present a significant challenge: while the extracted imagery is generally of sufficient quality and resolution, the associated captions are inadequate for RS image analysis. They are dominated by generic attributes and lack the detailed, domain-specific descriptions, thus fail to describe land cover types, specific geographical features, or other essential elements that are crucial in RS downstream tasks.
@@ II-B Remote Sensing Image-Text Paired Datasets
The increasing availability of RS images, coupled with the remarkable advancements in deep learning (DL), have generated unprecedented opportunities for automated satellite scene understanding, object detection, and change analysis at a global scale. However, the success of these data-intensive DL models hinges critically on the existence of large-scale, high-quality training datasets. While the computer vision community has witnessed the creation of massive image-text paired datasets like LAION-5B, the RS domain lags behind, primarily due to the inherent complexities and unique challenges associated with annotating RS imagery. Creating such datasets for RS applications (see Table I) is a formidable task, requiring domain expertise, meticulous manual effort, and substantial resources to accurately describe the complex spatial relationships, diverse land cover types, and varying scales present in aerial and satellite imagery.
@@
Building upon the aforementioned foundational datasets, Lu et al. [8] introduced the RSICD dataset, significantly expanding the scale and scope of RS image-text paired datasets available at the time. RSICD contains 10,921 aerial images featuring varying spatial resolutions, collected from multiple sources including Google Earth, Baidu Map, MapABC, and Tianditu, with the objective of achieving high intra-class variance and low inter-class similarity. The images were initially annotated with 24,333 descriptive sentences, with captions per image varying from 1 to 5 and subsequently extended them to 54,605 sentences by duplicating randomly the existing sentences when there are not five different sentences to describe the same image. While RSICD represented a significant advancement in terms of size and diversity compared to UCM-Captions and Sydney-Captions, it still relied on manual annotation, which limited its scalability and required substantial human effort.
@@
To address the need for fine-grained understanding of RS imagery, Yuan et al. [9] developed the RSITMD (Remote Sensing Image-Text Match) dataset specifically designed for cross-modal RS image retrieval. RSITMD comprises of 4,743 images across 32 categories, sourced from the RSICD dataset and Google Earth. Uniquely, RSITMD emphasizes fine-grained relationships between image attributes. Multiple annotators provided five descriptive captions per image, following guidelines to document these fine-grained relationships, alongside one to five supplementary object-level keywords. Authors demonstrated that RSITMD’s captions are more diverse than previous datasets, with a diversity score of 4.60, significantly higher than Sydney (1.83), UCM (0.97), and RSICD (1.67), capturing nuanced relationships between objects within images.
@@
Further expanding the scale and diversity of RS image captioning datasets, the NWPU-Captions dataset, introduced by Lu et al. [29], became the largest dataset for this task at the time. It contains 31,500 images with spatial resolution ranging from 30 to 0.2 meters, covering 45 classes based on the NWPU-RESISC45 dataset [30]. Each image was manually annotated with five descriptive captions by experienced RS experts, summing up to 157,500 image-caption pairs.
@@
The TextRS dataset, introduced by Abdullah et al. [31], focused on text-to-image matching tasks within the RS domain. TextRS contains approximately 2,144 images, each associated with five descriptive sentences summing up to 10,720 image-caption pairs. The images were randomly selected from four well-established scene classification datasets: AID [32], UC Merced [28], PatternNet [33], and NWPU [29], covering in total 134 overlapping categories. The annotation process involved five individuals generating descriptive sentences for each image, following specific guidelines to ensure diversity, accuracy, and relevance. While TextRS addressed limitations of previous datasets by incorporating diverse imagery, its reliance on manual annotation highlighted the ongoing need for automated or semi-automated approaches.
@@
The RSVGD dataset, introduced by Li et al. [10], attempted to address the annotation bottleneck by proposing an automatic image-query generation method, although it still required manual verification to ensure correctness. The dataset was sampled from the DIOR [34] dataset, originally designed for object detection, featuring 20 object categories. Authors annotation pipeline comprised of four steps: (1) box sampling, (2) attribute extraction, (3) expression generation, and 4) manual verification. The constructed dataset contains 17,402 RS images with spatial resolution ranging from 0.5m to 30m and 38,320 image-caption pairs, where each object instance in the RS image corresponds to a unique caption. The average caption length is 7.47 words and the vocabulary size is 100 words.
@@
Bountos et al. [35] introduced Hephaestus, a dataset designed for InSAR image analysis, offering a rich collection of manually annotated Sentinel-1 interferograms specifically designed for diverse computer vision tasks, including image captioning. The dataset has been manually annotated by a team of InSAR experts and contains 19,919 individual interferograms over 44 volcanoes across the globe, with each interferogram paired with a single, yet detailed caption. These captions provide comprehensive textual descriptions of each interferogram, including observed ground deformation, atmospheric contributions, and overall image quality. The InSAR images were acquired from the Comet-LiCS [36] portal, an automated system that processes Sentinel-1 data to generate interferograms.
@@
Hu et al. [45] introduced RSI-CAP and RSIEval, two interconnected datasets designed for training and evaluating vision-language models in RS. RSICap consists of 2,585 high-quality, human-annotated captions. The dataset included imagery derived from the DOTA [46] object detection dataset, featuring different satellite and aerial sensors such as GF-2, JL-1 and Google Earth, encompassing both color and panchromatic images with a wide range of spatial resolutions. The annotation process followed specific guidelines, including describing image attributes, object attributes, and overall scene, as well as including reasonable speculations based on visual content. RSICap captions have an average length of 60 words. RSIEval, on the other hand, was designed for benchmarking VLMs on RS image captioning (RSIC) and visual question answering (RSVQA) tasks. It contains 100 high-quality image-caption pairs (following RSICap annotation guidelines) and 936 diverse image-question-answer triplets, averaging nine questions per image. The questions covered four categories: object-related, image-related, scene-related, and reasoning-related. While RSI-CAP and RSIEval provided valuable resources for training and evaluation, their reliance on manual annotation limited their scalability.
@@
Muhtar et al. [48] introduced two new datasets: LHRS-Align and LHRS-Instruct. LHRS-Align is a large-scale dataset of 1.15 million RS image-caption pairs. Images, sourced from Google Earth at 1m resolution, were paired with geographic features from OpenStreetMap (OSM). The dataset’s annotation pipeline involved geo-aligning images with OSM features, pruning irrelevant attributes, and generating captions using Vicuna-v1.5-13B. LHRS-Instruct is a multi-modal instruction-following dataset for RS image understanding, including diverse tasks and complex visual reasoning data generated by GPT-4. The dataset was created by forming instruction datasets from public RS caption datasets (RSITMD [9] and NWPU [29]) through data cleaning and conversation generation with Vicuna-v1.5-13B, and selecting 15,000 LHRS-Align images, calculating bounding boxes from OSM data, and using GPT-4 to create complex instruction data. While LHRS datasets offer a high degree of automation and addressed diverse tasks, the reliance on specific language models and the manual construction of the benchmark dataset highlighted ongoing challenges.
Finally, Liu et al. [49] introduced Git-10M, a large-scale RS image-text dataset containing 10 million pairs, sourced by 30% from public datasets (Million-AID [41], GeoPile [50], SSL4EO-S12 [51], SkySript [16]) and 70% from Google Earth imagery. The dataset spans diverse geographical regions globally and includes various spatial resolutions ranging from 0.5m to 128m/pixel and geospatial metadata. They authors employ an automated annotation pipeline using GPT-4o, with prompts incorporating included geospatial metadata to generate semantically accurate descriptions.
@@
Dataset Year Images Image-Text Pairs Resolution (m/pixel) Source(s) Annotation Method
UCM-Captions [7] 2016 2,100 10,500 0.3 USGS National Map Urban Area Imagery Manual
Sydney-Captions [7] 2016 613 3,065 0.5 Google Earth Manual
RSICD [8] 2017 10,921 54,605 Variable Google Earth, Baidu Map,
MapABC, Tianditu
Manual
TextRS [31] 2020 2,144 10,720 Variable AID, UC Merced, PatternNet, NWPU Manual
RSITMD [9] 2022 4,743 23,715 Variable RSICD, Google Earth Manual
NWPU-Captions [29] 2022 31,500 157,500 30 – 0.2 NWPU-RESISC45 Manual
RSVGD [10] 2023 17,402 38,320 0.5 – 30 DIOR Hybrid
Hephaestus [35] 2023 19,919 19,919 100 Comet-LiCS Manual
RSI-CAP [45] 2023 2,585 2,585 Variable DOTA, GF-2, JL-1, Google Earth Manual
RSTeller [37] 2024 1.2M 2.5M 0.6 NAIP via Google Earth Engine Automatic
RS5M [17] 2024 5M 5M Variable PUB11 filtered datasets:
LAION-400M, YFCC100M
CC3M, CC12M
RedCaps, WIT
TextCaps, VizWiz
COCO, SBU
Conceptual Captions
RS3 datasets:
BigEarthNet-v1
FMoW
MillionAID
@@
SkyScript [16] 2024 2.6M 2.6M 0.1 – 30 Google Earth Engine Automatic
LuojiaHOG [43] 2024 94,856 94,856 Variable Google Earth Hybrid
ChatEarthNet [47] 2024 173,488 173,488 10 Sentinel-2, ESA WorldCover Automatic
LHRS-Align [48] 2024 1.15M 1.15M 1 Google Earth, OSM Automatic
Git-10M [49] 2025 10M 10M 0.5 – 128 30% public datasets (Million-AID, GeoPile, SSL4EO-S12, SkyScript), 70% Google Earth Automatic
GAIA (ours) 2025 41,030 205,150 Variable Several Remote Sensing-related websites Automatic

### v2

@@ Abstract
The continuous operation of Earth-orbiting satellites generates vast and ever-growing archives of Remote Sensing (RS) images. Natural language presents an intuitive interface for accessing, querying, and interpreting the data from such archives. However, existing Vision-Language Models (VLMs) are predominantly trained on web-scraped, noisy image-text data, exhibiting limited exposure to the specialized domain of RS. This deficiency results in poor performance on RS-specific tasks, as commonly used datasets often lack detailed, scientifically accurate textual descriptions and instead emphasize solely on attributes like date and location. To bridge this critical gap, we introduce GAIA, a novel dataset designed for multi-scale, multi-sensor, and multi-modal RS image analysis. GAIA comprises of 201,005 meticulously curated RS image-text pairs, representing a diverse range of RS modalities associated to different spatial resolutions. Unlike existing vision-language datasets in RS, GAIA specifically focuses on capturing a diverse range of RS applications, providing unique information about environmental changes, natural disasters, and various other dynamic phenomena. The dataset provides a spatially and temporally balanced distribution, spanning across the globe, covering the last 25 years with a balanced temporal distribution of observations. GAIA’s construction involved a two-stage process: (1) targeted web-scraping of images and accompanying text from reputable RS-related sources, and (2) generation of five high-quality, scientifically grounded synthetic captions for each image using carefully crafted prompts that leverage the advanced vision-language capabilities of GPT-4o. We also release an automated processing framework developed for this purpose, enabling the broader research community to generate captions for RS images using the web-crawled image-text data. Our extensive experiments, including fine-tuning of CLIP and BLIP2 models, demonstrate that GAIA significantly improves performance on RS image classification, cross-modal retrieval and image captioning tasks, proving its value as a crucial resource for advancing the field. We make our dataset, automated processing framework and fine-tuned model weights publicly available on our project’s GitHub repository [^1] https://github.com/Orion-AI-Lab/GAIA.
@@ I Introduction
The advancements in VLMs have drawn significant attention in the RS field, leading to studies across different applications. While early attempts have shown the potential of applying VLMs to RS, it is still an emerging field with many unsolved challenges [42]. A challenging issue that impedes the development of VLMs in RS is the lack of large-scale aligned image-text datasets. Existing RSI datasets mostly focus on visual recognition tasks and do not provide language annotations. While there have been some efforts to create RS image-text paired datasets, their scale, quality, and diversity still leave room for improvement.
@@
To address these challenges, we release GAIA, a dataset of approximately 40,201 RS images, each accompanied by five diverse and informative captions, resulting in 201,005 image-text pairs. Specifically, less that 3% of GAIA’s images are found in compared web-crawled datasets. Moreover, we provide extensive benchmarks for RS image classification and cross-modal retrieval tasks by fine-tuning CLIP on the GAIA dataset, and we publish the pre-trained model weights to facilitate further research. Complementing the dataset, we introduce a novel automated processing framework for generating RS image descriptions by leveraging web-crawled data from reputable sources and the advanced language generation capabilities of ChatGPT. Our work aims to bridge the gap between the vast potential of VLMs and the unique challenges of the RS domain, accelerating the development of foundation models that can unlock new insights from Earth observation data. We make our dataset, automated processing framework and fine-tuned model weights publicly available on our project’s GitHub repository [^3] https://github.com/Orion-AI-Lab/GAIA.
@@
    We introduce GAIA, a carefully curated RS dataset with 40,201 images along 5 synthetic captions each (201,005 image-text pairs) featuring global, multi-modal, multi-scale coverage spanning across the last 25 years. It exhibits minimal overlap ($3\%$) with existing datasets, enabling diverse Earth event analysis across multiple spatial resolutions and modalities.
@@ II Related work
VLMs have emerged as powerful tools for understanding and generating connections between visual and textual information. Pioneering works like CLIP [68] demonstrated the effectiveness of contrastive learning in creating aligned visual-semantic embeddings, while more recent models such as LLaVA [53, 52] have expanded these capabilities to include sophisticated visual reasoning and open-ended dialogue. Modern VLMs can perform a wide range of tasks, including image captioning, visual question answering, and cross-modal retrieval without task-specific fine-tuning. The success of these models stems from their ability to learn generalizable visual-semantic representations through self-supervised learning on vast amounts of image-text data, typically utilizing web-scale datasets containing billions of image-text pairs. This approach has proven particularly effective at capturing nuanced relationships between visual content and natural language descriptions, enabling zero-shot transfer to novel tasks and domains.
@@ II-A Have existing VLMs seen Remote Sensing imagery?
Several research efforts have focused on extracting RS subsets from web-crawled large-scale general-purpose datasets, extensively used within the computer vision domain. As an example, Czerkawski and Francis developed LAION-EO [19], leveraging CloudSEN12 [3] as an ”anchor dataset” of Sentinel-2 images to retrieve similar images from LAION-5B. Their multi-stage filtering process, involving nearest neighbor search and refined CLIP-based similarity thresholds, yielded 24k samples in the prototype version and an expanded set of 113k in version 1. Similarly, LAION-RS [84] identified a subset of 726k RS images, i.e. a mere 0.03% of LAION-2B, using a binary classifier. Zhang et. al. [97] introduced RS5M, a dataset of 5 million RS image-text pairs. A core component of RS5M, the PUB11 subset containing over 3 million pairs, was created by filtering 11 large-scale public datasets using keyword-based filtering, de-duplication, and a VLM-based filter. The 11 filtered large-scale public datasets, include LAION2B-en [70], LAION400M [73], LAIONCOCO [71], COYO700M [10], CC3M [75], CC12M [12], YFCC15M [81], WIT [76], Redcaps [20], SBU [63], and Visual Genome [36]. These filtering efforts offer critical insights into the current state of VLMs. First, the relatively small size of the extracted subsets, such as the minuscule 0.03% of LAION-2B forming LAION-RS, demonstrates the limited and incidental exposure of VLMs to RS data. Second, as shown in Fig. 1, web-scraped datasets like LAION-EO present a significant challenge: while the extracted imagery is generally of sufficient quality and resolution, the associated captions are inadequate for RS image analysis. They are dominated by generic attributes and lack the detailed, domain-specific descriptions, thus fail to describe land cover types, specific geographical features, or other essential elements that are crucial in RS downstream tasks.
@@ II-B Remote Sensing Image-Text Paired Datasets
The increasing availability of RS images, coupled with the remarkable advancements in DL, have generated unprecedented opportunities for automated satellite scene understanding, object detection, and change analysis at a global scale. However, the success of these data-intensive DL models hinges critically on the existence of large-scale, high-quality training datasets. While the computer vision community has witnessed the creation of massive image-text paired datasets like LAION-5B, the RS domain lags behind, primarily due to the inherent complexities and unique challenges associated with annotating RS imagery. Creating such datasets for RS applications (see Table I and Fig. 2) is a formidable task, requiring domain expertise, meticulous manual effort, and substantial resources to accurately describe the complex spatial relationships, diverse land cover types, and varying scales present in aerial and satellite imagery.
@@
Building upon the aforementioned foundational datasets, the RSICD dataset [57] was introduced, significantly expanding the scale and scope of RS image-text paired datasets available at the time. RSICD contains 10,921 aerial images featuring varying spatial resolutions, collected from multiple sources including Google Earth, Baidu Map, MapABC, and Tianditu, with the objective of achieving high intra-class variance and low inter-class similarity. The images were initially annotated with 24,333 descriptive sentences, with captions per image varying from 1 to 5 and subsequently extended them to 54,605 sentences by duplicating randomly the existing sentences when there are not five different sentences to describe the same image. While RSICD represented a significant advancement in terms of size and diversity compared to UCM-Captions and Sydney-Captions, it still relied on manual annotation, which limited its scalability and required substantial human effort.
@@
To address the need for fine-grained understanding of RS imagery, the RSITMD (Remote Sensing Image-Text Match) dataset [93] was specifically designed for cross-modal RS image retrieval. RSITMD comprises of 4,743 images across 32 categories, sourced from the RSICD dataset and Google Earth. Uniquely, RSITMD emphasizes fine-grained relationships between image attributes. Multiple annotators provided five descriptive captions per image, following guidelines to document these fine-grained relationships, alongside one to five supplementary object-level keywords. Authors demonstrated that RSITMD’s captions are more diverse than previous datasets, with a diversity score of 4.60, significantly higher than Sydney (1.83), UCM (0.97), and RSICD (1.67), capturing nuanced relationships between objects within images.
@@
Further expanding the scale and diversity of RS image captioning datasets, the NWPU-Captions dataset [16], became the largest dataset for this task at the time. It contains 31,500 images with spatial resolution ranging from 30 to 0.2 meters, covering 45 classes based on the NWPU-RESISC45 dataset [15]. Each image was manually annotated with five descriptive captions by experienced RS experts, summing up to 157,500 image-caption pairs.
@@
The TextRS dataset [1], focused on text-to-image matching tasks within the RS domain. TextRS contains approximately 2,144 images, each associated with five descriptive sentences summing up to 10,720 image-caption pairs. The images were randomly selected from four well-established scene classification datasets: AID [86], UC Merced [90], PatternNet [99], and NWPU [16], covering in total 134 overlapping categories. The annotation process involved five individuals generating descriptive sentences for each image, following specific guidelines to ensure diversity, accuracy, and relevance. While TextRS addressed limitations of previous datasets by incorporating diverse imagery, its reliance on manual annotation highlighted the ongoing need for automated or semi-automated approaches.
@@
The RSVGD dataset [96], attempted to address the annotation bottleneck by proposing an automatic image-query generation method, although it still required manual verification to ensure correctness. The dataset was sampled from the DIOR [40] dataset, originally designed for object detection, featuring 20 object categories. Authors annotation pipeline comprised of four steps: (1) box sampling, (2) attribute extraction, (3) expression generation, and 4) manual verification. The constructed dataset contains 17,402 RS images with spatial resolution ranging from 0.5m to 30m and 38,320 image-caption pairs, where each object instance in the RS image corresponds to a unique caption. The average caption length is 7.47 words and the vocabulary size is 100 words.
@@
Hephaestus [8], a dataset designed for InSAR image analysis, offers a rich collection of manually annotated Sentinel-1 interferograms specifically designed for diverse computer vision tasks, including image captioning. The dataset has been manually annotated by a team of InSAR experts and contains 19,919 individual interferograms over 44 volcanoes across the globe, with each interferogram paired with a single, yet detailed caption. These captions provide comprehensive textual descriptions of each interferogram, including observed ground deformation, atmospheric contributions, and overall image quality. The InSAR images were acquired from the Comet-LiCS [37] portal, an automated system that processes Sentinel-1 data to generate interferograms.
@@
RSICap and RSIEval [27], two interconnected datasets were designed for training and evaluating vision-language models in RS. RSICap consists of 2,585 high-quality, human-annotated captions. The dataset included imagery derived from the DOTA [85] object detection dataset, featuring different satellite and aerial sensors such as GF-2, JL-1 and Google Earth, encompassing both color and panchromatic images with a wide range of spatial resolutions. The annotation process followed specific guidelines, including describing image attributes, object attributes, and overall scene, as well as including reasonable speculations based on visual content. RSICap captions have an average length of 60 words. RSIEval, on the other hand, was designed for benchmarking VLMs on RS image captioning (RSIC) and visual question answering (RSVQA) tasks. It contains 100 high-quality image-caption pairs (following RSICap annotation guidelines) and 936 diverse image-question-answer triplets, averaging nine questions per image. The questions covered four categories: object-related, image-related, scene-related, and reasoning-related. While RSICap and RSIEval provided valuable resources for training and evaluation, their reliance on manual annotation limited their scalability.
@@
LHRS-Align [60] is a large-scale dataset of 1.15 million RS image-caption pairs. Images, sourced from Google Earth at 1m resolution, were paired with geographic features from OpenStreetMap (OSM). The dataset’s annotation pipeline involved geo-aligning images with OSM features, pruning irrelevant attributes, and generating captions using Vicuna-v1.5-13B. While the dataset offers a high degree of automation and addresses diverse tasks, the reliance on the effectiveness of a single language model highlights ongoing challenges.
Finally, Git-10M [50], a large-scale RS image-text paired dataset contains 10 million pairs, sourced by 30% from public datasets (Million-AID [55], GeoPile [59], SSL4EO-S12 [83], SkySript [84]) and 70% from Google Earth imagery. The dataset spans diverse geographical regions globally and includes various spatial resolutions ranging from 0.5m to 128m/pixel and geospatial metadata. They authors employ an automated annotation pipeline using GPT-4o, with prompts incorporating included geospatial metadata to generate semantically accurate descriptions.
@@
Dataset Year Images Image-Text Pairs Spatial Resolution (m/pixel) Spatial Coverage Temporal Coverage Image Modalities Observation Type Data Source(s) Annotation Method
Standard
Event-Triggered
UCM-Captions [67] 2016 2,100 10,500 0.3 USA - Optical (RGB) $\times$ USGS National Map Urban Area Imagery Manual
Sydney-Captions [67] 2016 613 3,065 0.5 Sydney, Australia - Optical (RGB) $\times$ Google Earth Manual
RSICD [57] 2017 10,921 54,605 Variable - - Optical (RGB) $\times$ Google Earth, Baidu Map, MapABC, Tianditu Manual
TextRS [1] 2020 2,144 10,720 Variable - - Optical (RGB) $\times$ AID, UC Merced, PatternNet, NWPU Manual
RSITMD [93] 2022 4,743 23,715 Variable - - Optical (RGB) $\times$ RSICD, Google Earth Manual
NWPU-Captions [16] 2022 31,500 157,500 30 – 0.2 - - Optical (RGB) $\times$ NWPU-RESISC45 Manual
RSVGD [96] 2023 17,402 38,320 0.5 – 30 - - Optical (RGB) $\times$ DIOR Hybrid
Hephaestus [8] 2023 19,919 19,919 100 44 most active volcanoes globally 2014-2021 Radar (InSAR) Comet-LiCS Manual
RSICap [27] 2023 2,585 2,585 Variable - - Optical (RGB) $\times$ DOTA, GF-2, JL-1, Google Earth Manual
RSTeller [25] 2024 1.2M 2.5M 0.6 USA 2021-2022 Optical (RGB) $\times$ NAIP via Google Earth Engine Automatic
RS5M [97] 2024 5M 5M Variable Global - Optical (RGB) $\times$ PUB11 filtered datasets: LAION-400M, YFCC100M, CC3M, CC12M, RedCaps, WIT, TextCaps, VizWiz, COCO, SBU, Conceptual Captions
RS3 datasets: BigEarthNet-v1, FMoW, MillionAID
@@
SkyScript [84] 2024 2.6M 2.6M 0.1 – 30 Global - Optical (RGB) $\times$ Google Earth Engine Automatic
LuojiaHOG [98] 2024 94,856 94,856 Variable Global - Optical (RGB) $\times$ Google Earth Hybrid
ChatEarthNet [92] 2024 173,488 173,488 10 Global 2020 Optical (Multi-spectral) $\times$ Sentinel-2, ESA WorldCover Automatic
LHRS-Align [60] 2024 1.15M 1.15M 1 Global - Optical (RGB) $\times$ Google Earth, OSM Automatic
Git-10M [50] 2025 10M 10M 0.5 – 128 Global - Optical (RGB) $\times$ 30% public datasets (Million-AID, GeoPile, SSL4EO-S12, SkyScript), 70% Google Earth Automatic
GAIA (ours) 2025 40,201 201,005 Variable Global 1998-2024 Multi-modal Visual Composites Web-scraped articles (images & text) from several RS-related websites Automatic
Figure 2: A qualitative comparison of remote sensing (RS) image-text datasets, highlighting differences in caption length, number of captions per RS image, level of detail, and domain specificity. Notably, our GAIA dataset (bottom right) features interpretative, context-aware, and RS domain-specific captions, which differs from the predominantly object-level descriptions of datasets like UCM-Captions and RSICD.